A compound containing chromium and sulfur contains 76.4 mass percent chromium.

Determine its empirical formula.


1
Expert's answer
2021-06-11T01:58:03-0400

Divisiblity factor for Cr "= \dfrac{weight percent}{molar mass } =\dfrac{76.4}{52}=1.47"



Divisiblity factor for S "=" "\dfrac{weight percent}{molar mass } =\dfrac{23.6}{32}=0.73"


Now divide the above both terms with the smallest number between the two so and that will be the new divisiblity factor of the both.


So new factor for Cr = 2

So new factor for S = 1


Hence the emperical formula will became = "Cr_2 S"